The function


may be used to model radioactive decay. Q represents the quantity remaining after t years; k is the decay constant. The decay constant for plutonium-240 is k = 0.00011. What is the half-life, in years?
1
Expert's answer
2017-10-25T15:34:07-0400
The half-life may be expressed by the following equation
τ=ln⁡2/k=0.693/k
The decay constant for plutonium-240 is k = 0.00011
Finally, we get
τ=0.693/0.00011=6300 years